# Arnold-Thom conjecture for the arrival time of surfaces

### v1

@@ 4.1. Uniqueness of limit flows
A closely related notion is the entropy introduced by Colding-Minicozzi [CM12]. For an $n$-dimensional hypersurface $\Sigma\subseteq\mathbb{R}^{n+1},$ its entropy is defined by
@@
By Huisken’s monotonicity [H90], given a mean convex flow666 A mean convex level set flow supports a Brakke flow, for which Huisken’s monotonicity applies. $M_{t}$ and a point $(x_{0},t_{0}),$ the density $\Theta_{M_{t}}(x_{0},t_{0},\cdot)$ and the entropy $\lambda\left(M_{(\cdot)}\right)$ are always monotone. Thus, the definitions are both valid as long as we start from an initial hypersurface with finite entropy. Since we always start from a smooth closed hypersurface, in particular, they are always well-defined.
@@
Let $M_{t}$ be a mean curvature flow and $(x_{0},t_{0})$ be a cylindrical singularity. We call $(x_{0},t_{0})$ a non-degenerate cylindrical singularity if all non-trivial special limit flows are modeled on multiplicity one cylinders.
Based on the definition, a priori, it is possible that at a non-degenerate cylindrical point, different limit flows are modeled on different types of cylinders (with axes of different dimensions and directions). In the following theorem, we prove that this cannot happen at a non-degenerate cylindrical point.
@@
Let $M_{t}$ be a mean curvature flow with finite entropy and $(x_{0},t_{0})$ is a singularity of the flow. If $(x_{0},t_{0})$ is a non-degenerate cylindrical singularity whose tangent flow is given by a $k$-cylinder, then all non-trivial special limit flows are modeled on $k$-cylinders with axes parallel to that of the tangent flow.
This leads to the first part of Theorem 1.4. We will use this to prove the Arnold-Thom conjecture for non-degenerated cylindrical singularities later.
@@
Let $(x_{i},t_{i})\rightarrow(0,0)$ be space-time points on $M_{t}$ and $M^{i}_{t}=\lambda_{i}\left(M_{\lambda_{i}^{-2}t+t_{i}}-x_{i}\right)$, where $\lambda_{i}\rightarrow\infty$. By Brakke’s compactness theorem [B78], after passing to a subsequence, $M^{i}_{t}$ converges to an integral Brakke flow $M_{t}^{\infty}$ ($t<0$). By the non-degenerate assumption, such a limit must be modeled on a cylinder in $\mathcal{C}_{\ell}$ for some $\ell\in\{0,1,...,n\}$. [^7] Note that this limit flow may not be centered at the origin. Moreover this limit flow is smooth for $t<0$, but may not be smooth at $t=0$. By Brakke’s regularity theorem [B78] (cf. [W05]), the convergence is smooth on compact sets of $\mathbb{R}^{n+1}\times(-\infty,0)$. By the lower semi-continuity given in Lemma 4.2, the entropy of any limit flow is no more than the entropy of the tangent flow, which is modeled on a $k$-cylinder. Therefore, $\ell\leq k$ or $\ell=n$ by Lemma 4.2. We aim to show that $\ell=k$ or $\ell=n$.
Suppose toward a contradiction that this does not hold. Then there would exist a limit flow $M_{t}^{\infty}$ as described above and modeled on an $\ell$-cylinder with $\ell<k.$
@@
For any $\varepsilon>0$, there exists $r_{0}=r_{0}(\varepsilon)$ with the following property. For any fixed $0<r_{1}<r_{2}\leq r_{0}$, $\frac{1}{\sqrt{-t}}M_{t}^{(x_{i},t_{i},r^{-1})}$ is $\varepsilon$-close to $\Gamma_{-1}$ for all sufficiently large $i$ whenever $t\in[-2,-\frac{1}{2}]$ and $r\in[r_{1},r_{2}]$.
@@
Recall that $\mathcal{C}^{0}_{k}$ is the space of all $k$-cylinders centered at the origin. Again, in both definitions, we don’t specify the directions of the cylinders. For each $m>1$, by Claim 4.5, we have $\rho_{i}\leq\frac{r_{0}}{m}$ for large $i$. Therefore, $\limsup\limits_{i\to\infty}\rho_{i}\leq\frac{r_{0}}{m}$. Letting $m\rightarrow\infty,$ we have
@@
In particular, $\rho_{i}<r_{0}$ for all large $i$.
From the geometry of shrinking cylinders, there exists $\mathcal{T}=\mathcal{T}(\varepsilon)\geq 1$ such that $\frac{1}{\sqrt{-t}}M_{t}^{\infty}$ is $\frac{\varepsilon}{2}$-close to an $\ell$-cylinder for all $t\leq-\mathcal{T}$. We emphasize again that there is no specification for the directions of axes $\mathbb{R}^{\ell}$ here. By the smooth convergence, the rescaled flow $\frac{1}{\sqrt{\mathcal{T}}}M_{-\mathcal{T}}^{\left(x_{i},t_{i},\lambda_{i}\right)}$ is $\varepsilon$-close to an $\ell$-cylinder for all large $i$. Since $\frac{1}{\sqrt{\mathcal{T}}}M_{-\mathcal{T}}^{\left(x_{i},t_{i},\lambda_{i}\right)}=M_{-1}^{\left(x_{i},t_{i},\lambda_{i}\sqrt{\mathcal{T}}^{-1}\right)},$ we have $\underline{\rho}_{i}\geq\lambda_{i}^{-1}\sqrt{\mathcal{T}}>0$ for all large $i$. On the other hand, the definitions of $\rho_{i}$ and $\underline{\rho}_{i}$ imply $\underline{\rho}_{i}<\rho_{i}.$ Combining these, we have
@@
Now we consider the rescaled flow $M_{t}^{(x_{i},t_{i},\underline{\rho}_{i}^{-1})}$. The fact that $\lim\limits_{i\to\infty}\frac{\rho_{i}}{\underline{\rho}_{i}}=\Lambda$ and the definition of $M_{t}^{a}$ implies that $M_{t}^{(x_{i},t_{i},\underline{\rho}_{i}^{-1})}$ converges to $\Lambda M_{\Lambda^{-2}t}^{a}$. On the other hand, the definition of $\underline{\rho}_{i}$ implies that the limit (which is $\Lambda M_{\Lambda^{-2}t}^{a}$) must be $2\varepsilon$-close to an $\ell$-cylinder in $\mathcal{C}^{0}_{\ell}$ at $t=-1$. In other words, some time slice of a shrinking $k$-cylinder is $2\varepsilon$-close to an $\ell$-cylinder in $C_{\ell}^{0}$. However, this can never happen by the geometry of cylinders, a contradiction.
@@
We first show that at scale $\rho_{i}$ and above, not just the flow looks close to a $k$-cylinder centered at origin, but also the axis of the cylinder almost aligns with the tangent flow.
@@
for $\tau>0.$ By the definition of $\rho_{i}$ and the claim, $\overline{M}_{\tau}^{i}$ is $\varepsilon$-close to a $k$-cylinder for all $\tau\in\left[0,2\log\left(\frac{r_{0}}{\rho_{i}}\right)+\ln 2\right]$. By the entropy bound and Colding-Minicozzi’s Łojasiewicz inequality [CM15], the axis of the cylinder has total rotation at most $C\varepsilon$ during $\tau\in\left[1,2\log\left(\frac{r_{0}}{\rho_{i}}\right)\right]$ (c.f. [GH, Proposition 4.1] and [CM15, Lemma A.48]). By Claim 4.5, $\overline{M}_{\tau}^{i}$ is $\varepsilon$-close to $\Gamma_{-1}$ for $\tau\in[0,10]$ as long as $i$ is large enough. Therefore, $\overline{M}_{\tau}^{i}$ is $C\varepsilon$-close to $\Gamma_{-1}$ for all $\tau\in\left[0,2\log\left(\frac{r_{0}}{\rho_{i}}\right)\right]$. In particular, for all large $i,$ we have
@@
Before we end this section, we note that in [SX4], Sun-Xue defined a different notion of non-degenerate cylindrical singularities and proved that they are generic. Their non-degeneracy condition implies the one defined in Definition 4.3. Also, there are higher dimensional dumbbells having non-degenerate cylindrical singularities.
@@ 4.2. Uniqueness of limit points
The key idea is that Theorem 4.4 guarantees the existence of a “canonical neighborhood,” inside which every regular point behaves nicely, that is, like one on a cylindrical flow. This prevents the occurrence of multiple limit points.
@@
We assume the origin is a limit point of $\gamma$ and is a singularity of the flow modeled on a $k$-cylinder. With a reparametrization (using the time of the flow) and a translation, we assume $\gamma(t_{i})\to 0$ for some $t_{i}\to 0.$ Without loss of generality, we may assume that $\{t_{i}\}$ is a strictly increasing sequence.
For $i\geq 2$, we define
(4.14) $\delta_{i}=2|\gamma(t_{i})|+3\sqrt{n|t_{i-1}|}$
and
(4.15) $s_{i}=\inf\left\{\tau\leq t_{i}:(\gamma(t),t)\in B_{\delta_{i}}(0)\times\left[-\delta_{i}^{2},0\right)\text{ for all }t\in[\tau,t_{i}]\right\}.$
Clearly, $s_{i}<t_{i}<0$ and $\delta_{i}\rightarrow 0$. Moreover, the definition implies that $\left(\gamma(s_{i}),s_{i}\right)$ must be on the boundary of the parabolic neighborhood, which means
(4.16) $|\gamma(s_{i})|=\delta_{i}\text{ or }s_{i}=-\delta_{i}^{2}.$
@@
For large $i$ and $t\in[s_{i},t_{i}],$ $H\left(\gamma(t),t)\right)>0$ and
(4.18) $\lim\limits_{i\rightarrow\infty}\sup\limits_{t\in[s_{i},t_{i}]}\left|\frac{\partial_{t}H}{H^{3}}\left(\gamma(t),t)\right)-\frac{1}{n-k}\right|=0.$
@@
If the claim failed, then there would exist $\varepsilon>0,$ a subsequence of $i$ (still denoted by $i$), and $\tau_{i}\in[s_{i},t_{i}]$ such that
@@
(4.20) $H(\gamma(\tau_{i}),\tau_{i})=0.$
To get a contradiction, we consider
@@
By the definition of $\tau_{i},$ we also have $\gamma(\tau_{i})\to 0,$ which is a limit point, so $H(\gamma(\tau_{i}),\tau_{i})\rightarrow\infty$ because $\nabla u(0)=0.$ By Brakke’s compactness theorem [B78] (cf. [HK1]), after passing to a subsequence (still denoted by $i$), $M_{t}^{i}$ converges to an integral Brakke flow $M_{t}^{\infty}(t<0)$ By the non-degeneracy assumption and Theorem 4.4, $M_{t}^{\infty}$ is a shrinking cylinder modeled on a $k$-cylinder with $H(0,0)=1$. By the evolution of cylinders and the curvature control, we know that $\partial_{t}H=\frac{H^{3}}{n-k}$ on $M_{t}^{\infty}$ for $t<\sqrt{\frac{n-k}{2}}$. Brakke’s regularity theorem [B78] (cf. [W05, HK1]) guarantees that the convergence is smooth on compact sets in $\mathbb{R}^{n+1}\times(-\infty,0]$. Thus, for all large $i,$ we have
$\left|\frac{\partial_{t}H}{H^{3}}\left(\gamma(\tau_{i}),\tau_{i}\right)-\frac{1}{n-k}\right|<\frac{\varepsilon}{2},\text{ and }H\left(\gamma(\tau_{i}),\tau_{i}\right)>0.$
@@
We continue the proof of Theorem 4.13. By Claim 4.17, we have
@@
when $t\in[s_{i},t_{i}]$ for large $i.$ This differential inequality and the fact that $H(\gamma(t_{i}),t_{i})>0$ imply
$\frac{d}{dt}H(\gamma(t),t)^{-2}\leq-\frac{2}{n}.$
Integrating from $t$ to $t_{i},$ we obtain that
@@
From this, we claim that
(4.22) $s_{i}<-\frac{\delta_{i}^{2}}{9n}.$
Indeed, by (4.16), we have $\left|\gamma(s_{i})\right|=\delta_{i}$ or $s_{i}=-\delta_{i}^{2}.$ If $s_{i}=-\delta_{i}^{2},$ then (4.22) is automatically true. Otherwise, we have $\left|\gamma(s_{i})\right|=\delta_{i}.$ By the definition (4.14), we have $|\gamma(t_{i})|<\frac{\delta_{i}}{2}$. Combining these with (4.21), we get
$\sqrt{2n|s_{i}|}\geq|\gamma(s_{i})|-|\gamma(t_{i})|>\delta_{i}-\frac{\delta_{i}}{2}=\frac{\delta_{i}}{2},$
which implies $|s_{i}|\geq\frac{\delta_{i}^{2}}{8n},$ leading to (4.22) since $s_{i}<0.$
We are able to finish the proof. By (4.22), the definition (4.14) of $\delta_{i}$ immediately implies that $s_{i}<t_{i-1}$. This and the definition (4.15) of $s_{i}$ implies
$\gamma([t_{i-1},t_{i}])\subset B_{\delta_{i}}(0)$
for all large $i$. Since $\delta_{i}\to 0,$ for all fixed $\delta>0$, we have
$\gamma\left([t_{i-1},t_{i}]\right)\subset B_{\delta}(0)$
@@
$\gamma([t_{i},0))\subset B_{\delta}(0)$
for all large $i$. Because $\delta$ is arbitrary, we conclude that $\lim\limits_{t\rightarrow 0^{-}}\gamma(t)=0.$ This finished the proof of the uniqueness of limit points of $\gamma.$

### v2

@@ 4.1. Uniqueness of limit flows
A closely related notion is the entropy introduced by Colding–Minicozzi [CM12]. For an $n$-dimensional hypersurface $\Sigma\subseteq\mathbb{R}^{n+1},$ its entropy is defined by
@@
By Huisken’s monotonicity [H90], given a mean convex flow777 A mean convex level set flow supports a Brakke flow, for which Huisken’s monotonicity applies. $M_{t}$ and a point $(x_{0},t_{0}),$ the density $\Theta_{M_{t}}(x_{0},t_{0},\cdot)$ and the entropy $\lambda\left(M_{(\cdot)}\right)$ are always monotone. Thus, the definitions are both valid as long as we start from an initial hypersurface with finite entropy. Since we always start from a smooth closed hypersurface, in particular, they are always well-defined.
@@
Let $M_{t}$ be a mean curvature flow and $(x_{0},t_{0})$ be a cylindrical singularity. We call $(x_{0},t_{0})$ a non-degenerate cylindrical singularity if all non-trivial special limit flows are multiplicity one cylinders.
Based on the definition, a priori, it is possible that at a non-degenerate cylindrical point, different limit flows are different types of cylinders (with axes of different dimensions and directions). In the following theorem, we prove that this cannot happen at a non-degenerate cylindrical point.
@@
Let $M_{t}$ be a mean curvature flow with finite entropy and $(x_{0},t_{0})$ is a singularity of the flow. If $(x_{0},t_{0})$ is a non-degenerate cylindrical singularity whose tangent flow is given by a $k$-cylinder, then all non-trivial special limit flows are $k$-cylinders with axes parallel to that of the tangent flow.
This leads to the first part of Theorem 1.4. We will use this to prove the Arnold–Thom conjecture for non-degenerated cylindrical singularities later.
@@
Let $(x_{i},t_{i})\rightarrow(0,0)$ be space-time points on $M_{t}$ and $M^{i}_{t}=\lambda_{i}\left(M_{\lambda_{i}^{-2}t+t_{i}}-x_{i}\right)$, where $\lambda_{i}\rightarrow\infty$. By Brakke’s compactness theorem [B78], after passing to a subsequence, $M^{i}_{t}$ converges to an integral Brakke flow $M_{t}^{\infty}$ ($t<0$). By the non-degenerate assumption, such a limit must be a cylinder in $\mathcal{C}_{\ell}$ for some $\ell\in\{0,1,...,n\}$. [^8] Note that this limit flow may not be centered at the origin. Moreover this limit flow is smooth for $t<0$, but may not be smooth at $t=0$.
By Brakke’s regularity theorem [B78] (cf. [W05]), the convergence is smooth on compact sets of $\mathbb{R}^{n+1}\times(-\infty,0)$. By the lower semi-continuity given in Lemma 4.2, the entropy of any limit flow is no more than the entropy of the tangent flow, which is a $k$-cylinder. Therefore, $\ell\leq k$ or $\ell=n$ by Lemma 4.2. We aim to show that $\ell=k$ or $\ell=n$.
We remark that the argument below will also give us a contradiction if $\ell>k.$ Therefore, the entropy argument above is not completely necessary.
Suppose toward a contradiction that this does not hold. Then there would exist a limit flow $M_{t}^{\infty}$ as described above and equal to an $\ell$-cylinder with $\ell<k.$
@@
For any $\varepsilon>0$, there exists $r_{0}=r_{0}(\varepsilon)$ with the following property. For any fixed $0<r_{1}<r_{2}\leq r_{0}$, $\frac{1}{\sqrt{-t}}M_{t}^{(x_{i},t_{i},r^{-1})}$ is $\varepsilon$-close [^9] In the rest of Section 4, we say $M$ is $\varepsilon$-close to $N$ if $M$ can be written as the graph of a function over $N$ in $C^{\lceil 1/\varepsilon\rceil}\left(B_{1/\varepsilon}\right)$ with norm at most $\varepsilon.$ to $\Gamma_{-1}$ for all sufficiently large $i$ whenever $t\in[-2,-\frac{1}{2}]$ and $r\in[r_{1},r_{2}]$.
@@
Recall that $\mathcal{C}^{0}_{k}$ is the space of all $k$-cylinders centered at the origin. Again, in both definitions, we don’t specify the directions of the cylinders. For each $m>1$, by Claim 4.5, we have $\rho_{i}\leq\frac{r_{0}}{m}<r_{0}$ for large $i$. Therefore, $\limsup\limits_{i\to\infty}\rho_{i}\leq\frac{r_{0}}{m}$. Letting $m\rightarrow\infty,$ we have
@@
Since $M_{t}^{\infty}$ is a shrinking $\ell$-cylinder, its axis keeps a fixed distance to the origin. If we scale it by $\frac{1}{\sqrt{-t}}$ then the axis will get closer to the origin when $-t$ gets larger. As a result, there exists $\mathcal{T}=\mathcal{T}(\varepsilon)\geq 1$ such that $\frac{1}{\sqrt{-t}}M_{t}^{\infty}$ is $\frac{\varepsilon}{2}$-close to an $\ell$-cylinder for all $t\leq-\mathcal{T}$.We emphasize again that there is no specification for the directions of axes $\mathbb{R}^{\ell}$ here. By the smooth convergence, the rescaled flow $\frac{1}{\sqrt{\mathcal{T}}}M_{-\mathcal{T}}^{\left(x_{i},t_{i},\lambda_{i}\right)}$ is $\varepsilon$-close to an $\ell$-cylinder for all large $i$. Since $\frac{1}{\sqrt{\mathcal{T}}}M_{-\mathcal{T}}^{\left(x_{i},t_{i},\lambda_{i}\right)}=M_{-1}^{\left(x_{i},t_{i},\lambda_{i}\sqrt{\mathcal{T}}^{-1}\right)},$ we have $\underline{\rho}_{i}\geq\lambda_{i}^{-1}\sqrt{\mathcal{T}}>0$ for all large $i$. On the other hand, the definitions of $\rho_{i}$ and $\underline{\rho}_{i}$ imply $\underline{\rho}_{i}<\rho_{i}.$ Combining these, we have
@@
Now, we consider the rescaled flow $M_{t}^{(x_{i},t_{i},\underline{\rho}_{i}^{-1})}$. The fact that $\lim\limits_{i\to\infty}\frac{\rho_{i}}{\underline{\rho}_{i}}=\Lambda$ and the definition of $M_{t}^{a}$ implies that $M_{t}^{(x_{i},t_{i},\underline{\rho}_{i}^{-1})}$ converges to $\Lambda M_{\Lambda^{-2}t}^{a}$. On the other hand, the definition of $\underline{\rho}_{i}$ implies that the limit (which is $\Lambda M_{\Lambda^{-2}t}^{a}$) must be $2\varepsilon$-close to an $\ell$-cylinder in $\mathcal{C}^{0}_{\ell}$ at $t=-1$. In other words, some time slice of a shrinking $k$-cylinder is $2\varepsilon$-close to an $\ell$-cylinder in $C_{\ell}^{0}$. This cannot happen when $\varepsilon$ is small enough. (For example, we can take $\varepsilon<(100n)^{-1}$ to prevent this.)
@@
We first show that at scale $\rho_{i}$ and above, not only does the flow look close to a $k$-cylinder centered at origin, but also the axis of the cylinder almost aligns with the tangent flow.
@@
for $\tau>0.$ By the definition of $\rho_{i}$ and the claim, $\overline{M}_{\tau}^{i}$ is $\varepsilon$-close to a $k$-cylinder for all $\tau\in\left[0,2\log\left(\frac{r_{0}}{\rho_{i}}\right)+\log 2\right]$. By the entropy bound and Colding–Minicozzi’s Łojasiewicz inequality [CM15], the axis of the cylinder has total rotation at most $C\varepsilon$ during $\tau\in\left[1,2\log\left(\frac{r_{0}}{\rho_{i}}\right)\right]$ (cf. [GH, Proposition 4.1] and [CM15, Lemma A.48]). By Claim 4.5, $\overline{M}_{\tau}^{i}$ is $\varepsilon$-close to $\Gamma_{-1}$ for $\tau\in[0,10]$ as long as $i$ is large enough. Therefore, $\overline{M}_{\tau}^{i}$ is $C\varepsilon$-close to $\Gamma_{-1}$ for all $\tau\in\left[0,2\log\left(\frac{r_{0}}{\rho_{i}}\right)\right]$. In particular, for all large $i,$ we have
@@
Before we end this section, we note that in [SX4], Sun–Xue defined a different notion of non-degenerate cylindrical singularities and proved that they are generic. Their non-degeneracy condition implies the one defined in Definition 4.3. Also, there are higher dimensional dumbbells having non-degenerate cylindrical singularities. In general, examples of higher dimensional cylindrical singularities were studied in [W13].
@@ 4.2. Uniqueness of limit points
The key idea is that Theorem 4.4 guarantees the existence of a “canonical neighborhood,” inside which every regular point behaves nicely, that is, like one on a cylindrical flow. This prevents the occurrence of multiple limit points. As mentioned in Remark 4.12, since Theorem 4.4 does not require mean convexity, we can also extend Theorem 4.13 to a general flow with a non-degenerate cylindrical singularity if we use a general notion of flow lines, cf. Definition 5.2.
@@
We assume the origin is a limit point of $\gamma$ and is a singularity of the flow modeled on a $k$-cylinder. With a reparametrization (using the time of the flow) and a translation, we assume $\gamma(t_{i})\to 0$ for some $t_{i}\to 0.$ Our goal is to show $\gamma(t)\to 0$ as $t\to 0.$ We will prove this by showing that $\gamma$ behaves like a cylindrical flow lines for $t\in[s_{i},t_{i}]$ for some $s_{i}<t_{i},$ and use this cylindricality to show that we can in fact take $s_{i}=t_{i-1}.$ Without loss of generality, we may assume that $\{t_{i}\}$ is a strictly increasing sequence.
Fix any $\delta>0$, for $i\geq 2$, we let
(4.14) $s_{i}=s_{i}(\delta)=\inf\left\{\tau\in[t_{i-1},t_{i}]:|\gamma(t)|\leq|\gamma(t_{i})|+\delta\text{ for all }t\in[\tau,t_{i}]\right\}.$
@@
Given any $\varepsilon>0,$ there exists $\delta_{0}=\delta_{0}(\varepsilon)>0$ with the following property. If $\delta\leq\delta_{0}$, then for all large $i$ and $t\in[s_{i}(\delta),t_{i}],$ we have $H\left(\gamma(t),t\right)>0$ and
(4.16) $\sup\limits_{t\in[s_{i}(\delta),t_{i}]}\left|\frac{\partial_{t}H}{H^{3}}\left(\gamma(t),t\right)-\frac{1}{n-k}\right|\leq\varepsilon.$
@@
If the claim were false, then there would exist $\varepsilon>0,$ $\delta_{i}\rightarrow 0$, $j_{i}\geq i$ and $\tau_{i}\in[s_{j_{i}}(\delta_{i}),t_{j_{i}}]$ such that
@@
(4.18) $H(\gamma(\tau_{i}),\tau_{i})=0$
for each $i$. We may pass to a subsequence (still denoted by $i$) so that $j_{i}$ is strictly increasing in $i$. To get a contradiction, we consider
@@
By definition we have $\tau_{i}\rightarrow 0$. Moreover,
$|\gamma(\tau_{i})|\leq|\gamma(t_{j_{i}})|+\delta_{i}\rightarrow 0$
Thus $\gamma(\tau_{i})\rightarrow 0$, which is a limit point, so $H(\gamma(\tau_{i}),\tau_{i})\rightarrow\infty$ because $\nabla u(0)=0.$ By Brakke’s compactness theorem [B78] (cf. [HK1]), after passing to a subsequence (still denoted by $i$), $M_{t}^{i}$ converges to an integral Brakke flow $M_{t}^{\infty}.$ By the non-degeneracy assumption and Theorem 4.4, $M_{t}^{\infty}$ is a shrinking $k$-cylinder with $H(0,0)=1$. By the evolution of cylinders, we know that $\partial_{t}H=\frac{H^{3}}{n-k}$ on $M_{t}^{\infty}$ for $t<\sqrt{\frac{n-k}{2}}$. Brakke’s regularity theorem [B78] (cf. [W05, HK1]) guarantees that the convergence is smooth on compact sets in $\mathbb{R}^{n+1}\times(-\infty,0]$. Thus, for all large $i,$ we have
$\left|\frac{\partial_{t}H}{H^{3}}\left(\gamma(\tau_{i}),\tau_{i}\right)-\frac{1}{n-k}\right|\leq\varepsilon\text{ and }H\left(\gamma(\tau_{i}),\tau_{i}\right)>0.$
@@
We continue the proof of Theorem 4.13. We specifically choose $\varepsilon=1/n^{2}$ and take $\delta_{0}$ to be given by Claim 4.15. We then work with any fixed $\delta<\delta_{0}$ and write $s_{i}=s_{i}(\delta).$ Then we have
@@
when $t\in[s_{i},t_{i}]$ for all large $i.$ This differential inequality and the fact that $H(\gamma(t_{i}),t_{i})>0$ imply
$\frac{d}{dt}H(\gamma(t),t)^{-2}\leq-\frac{2}{n}$
when $t\in[s_{i},t_{i}].$ Integrating from $t$ to $t_{i},$ we obtain that
@@
where we use that for all large $i$, we have
$\sqrt{2n(t_{i}-s_{i})}\leq\sqrt{2n(t_{i}-t_{i-1})}<\delta.$
On the other hand, by the definition of $s_{i},$ if $s_{i}\neq t_{i-1},$ then $|\gamma(s_{i})|=|\gamma(t_{i})|+\delta.$ This contradicts (4.19). Thus, we deduce that $s_{i}=t_{i-1}$ for all large $i$. This then implies
$\gamma([t_{i-1},t_{i}])\subset B_{|\gamma(t_{i})|+\delta}(0)$
for all large $i$. Because $|\gamma(t_{i})|\to 0,$ we have
$\gamma\left([t_{i-1},t_{i}]\right)\subset B_{2\delta}(0)$
@@
$\gamma([t_{i},0))\subset B_{2\delta}(0)$
for all large $i$. Because $\delta<\delta_{0}$ can be arbitrarily small, we conclude that $\lim\limits_{t\rightarrow 0^{-}}\gamma(t)=0.$ This finished the proof of the uniqueness of limit points of $\gamma.$
Remark 4.20.
The proof of Theorem 4.13 may lead to another proof for Theorem 3.1 when the flow line is cylinder-type. Without knowing Theorem 3.1 a priori, one may still prove Theorems 3.7 and 3.9. The possibility of having more than one limit point may complicate the proof, requiring the Łojasiewicz estimate by Colding–Minicozzi in the argument again. After the uniqueness of limit flows is proven in this way, to prove the uniqueness of limit points, the case of cylinder-type flow lines can be treated as in the proof of Theorem 4.13, while the case of bowl-type flow lines can then be dealt with directly, as our treatment of Case 1 in the proof of Theorem 3.16.
Remark 4.21.
The techniques used to prove Claim 4.15 share similarities with those in Lemma 3.10 (cf. equation (3.11)), as both rely on the differential inequalities for $H$ derived from the Jacobi equation and approximation by a cylinder. A key difference, however, is that in Lemma 3.10, the uniqueness of limit points is already known, and equation (3.11) is used to establish the angle estimate (3.14). In contrast, in the proof of Theorem 4.13, the cylindrical estimate must be handled more delicately due to the absence of such uniqueness.